Question 11 (1 point) There are 100 people who got HIV tested. Among them 15% gt HIV disease and 를 of people who got HIV disease had HIV test positive Among 100 people, 45 people got HIlV test negative and HIV disease. Find the specificity (choose the closest answer) 0.42 0.52 0.32 。0.72 0.62 Save : Question 12 (1 point) There are 100 people who got HIV tested. Among them 15% got HIV disease and 를 of people who got HIV disease had HiV test positive. Among 100 people, 45 people got HIV test negative and HiV disease Find the sensitivity fchoose the closest answer 0.46 07e
Answer #1

The following is the small explanation about sensitivity and specificity.

   ACTUAL CONDITION
TEST RESULT

  

Positive

Disease ​ Non-Disease Total

A

(True Positives)

B

(False Positive)

T

(Test Positive)

Negative

   C

(False Negative)

D

(True Negative)

T

(Test Negative)

   T (Disease) T(Non-Disease) Total

Sensitivity is the probability that a test result will indicate 'disease' among those with the disease:

SENSITIVITY =

Specificity is the probability of those without disease who will have a negative test result:

SPECIFICITY =

QUESTION 11:

   ACTUAL CONDITION (HIV)
HIV TEST RESULT

  

Positive

Disease ​ Non-Disease Total

   10

(True Positives)

40

(False Positive)

50

(Test Positive)

Negative

5

(False Negative)

45

(True Negative)

50

(Test Negative)

15    85 100

Total number of people who got HIV tested = 100

Number of people who got HIV disease= 15% of 100=(15/100)*100=15

Number of people who got HIV disease and test result HIV positive (True Positives)=(2/3)*15=10

Number of people with test result negative and HIV disease present (False Negative)=15-10=5

Number of people without HIV disease and test result HIV negative (True Negatives)= 45

As we knew the total of HIV tested people is 100 and total of diseased people is 15, we can calculate the total of non-diseased people=100-15=85.

And as we calculated T( Non-diseased)=85 and True negatives is 45, False positives is calculated as T(Non-diseased)-True Negative=85-45=40.

SPECIFICITY =

%

Thus 54% of the people without HIV disease are correctly identified as negative by HIV test. (That is option (b) 0.52 is the closest answer).

QUESTION 12:

   ACTUAL CONDITION (HIV)
HIV TEST RESULT

  

Positive

Disease ​ Non-Disease Total

   10

(True Positives)

40

(False Positive)

50

(Test Positive)

Negative

5

(False Negative)

45

(True Negative)

50

(Test Negative)

15    85 100

Total number of people who got HIV tested = 100

Number of people who got HIV disease= 15% of 100=(15/100)*100=15

Number of people who got HIV disease and test result HIV positive (True Positives)=(2/3)*15=10

Number of people with test result negative and HIV disease present (False Negatives)=15-10=5

Number of people without HIV disease and test result HIV negative (True Negative)= 45

As we knew the total of HIV tested people is 100 and total of diseased people is 15, we can calculate the total of non-diseased people=100-15=85.

And as we calculated T( Non-diseased)=85 and True negatives is 45, False positives is calculated as T(Non-diseased)-True Negative=85-45=40.

SENSITIVITY =

%

Thus 67% of the people with HIV disease are correctly identified as positive by HIV test. (That is option (b) 0.66 is the closest answer)